The present ages of Andy and Mike are in the ratio 5 : 3. If Andy had been 7 years older and Mike 7 years younger, the age of Andy would have been 3 times the age of Mike. Find their present ages.

1
Expert's answer
2021-03-09T07:51:25-0500

Let Andy be "5x" years old and let Mike be "3x" years old.

If Andy had been 7 years older, the age of Andy would have been "5x+7".

If Mike had been 7 years younger, the age of Andy would have been "3x-7"

Also we know that age of Andy would have been 3 times the age of Mike:

"5x+7=3(3x-7)"

"5x+7=9x-21"

"28=4x"

"x=7"


"5x=5\cdot7=35" and "3x=3\cdot7=21"

Answer: Andy is 35 years old and Mike is 21 years old.